OpenStudy (anonymous):

The sum of two numbers is 64. Twice the larger number plus five times the smaller number is 20. Find the two numbers.

OpenStudy (anonymous):

Let x and y be the two numbers and x > y. From your story: $$x+y = 64;$$ $$2x +5y = 20$$ which gives $$ \{\{x\to 100,y\to -36\}\}$$

OpenStudy (anonymous):

You can use simple row reduction to work out these problems. a + b = 64 2a + 5b = 20 [1 1][64] [2 5][20] [1 1][64] [0 3][-108] R2 - 2R1 You already have your answer here (-108/3 = -36, 64--36 = 100) but I'll finish anyway: [1 1][64] [0 1][-36] R2/3 [1 0][100] R1 - R2 [0 1][-36] a = 100, b = -36

OpenStudy (anonymous):

^^ That is not the solution of someone who have few screw missing

OpenStudy (anonymous):

btw is this a kind of short cut for the using finding the inverse and then multiply ?

OpenStudy (anonymous):

Basically, yes. You can calculate the inverse of the product matrix and multiply it by the resultant matrix and have the same result. Or if you have a graphics calculator (or a calculator that supports matrices) it can do it for you. It's the first way I learnt to solve systems of linear equations. The inverse way takes a little longer. Since we have a + b = 64 and 2a + 5b = 20 we can assign it to a matrix: \[\Large \left[ {\begin{array}{*{20}{c}} a&b\\ {2a}&{5b} \end{array}} \right] = \left[ {\begin{array}{*{20}{c}} {64}\\ {20} \end{array}} \right]\] To work it out via an inverse matrix, just find the determinant of the square matrix: \[\Large \begin{array}{l} \left[ {\begin{array}{*{20}{c}} a&b\\ c&d \end{array}} \right] = ad - cb\\ \left[ {\begin{array}{*{20}{c}} 1&1\\ 2&5 \end{array}} \right] = 1 \cdot 5 - 2 \cdot 1 = 3 \end{array}\] Then flip the product matrix around and multiply it be the inverse of 3 (i.e. 1/3): \[\Large \left[ {\begin{array}{*{20}{c}} 5&{ - 1}\\ { - 2}&1 \end{array}} \right] \cdot \frac{1}{3} = \left[ {\begin{array}{*{20}{c}} {5/3}&{ - 1/3}\\ { - 2/3}&{1/3} \end{array}} \right]\] Then multiply that by the resultant matrix: \[\Large \left[ {\begin{array}{*{20}{c}} {5/3}&{ - 1/3}\\ { - 2/3}&{1/3} \end{array}} \right] \cdot \left[ {\begin{array}{*{20}{c}} {64}\\ {20} \end{array}} \right] = \left[ {\begin{array}{*{20}{c}} {(5/3) \cdot 64 + ( - 1/3) \cdot 20}\\ {( - 2/3) \cdot 64 + (1/3) \cdot 20} \end{array}} \right] = \left[ {\begin{array}{*{20}{c}} {100}\\ { - 36} \end{array}} \right]\]

OpenStudy (anonymous):

Well by flipping the product matrix around (s)he meant swap the position of the diagonal elements and change the sign of the off-diagonal.

OpenStudy (anonymous):

He ;) And yes, that's what I meant. I suppose I should have been clearer, 'flipping' sounds like simply turning upside down: \[\Large \left[ {\begin{array}{*{20}{c}} a&b\\ c&d \end{array}} \right] \Rightarrow \left[ {\begin{array}{*{20}{c}} d&{ - b}\\ { - c}&a \end{array}} \right]\] I haven't played with matrices since... a long time ago... so I can't really remember all the rules - but point of the story is, row reduction is handy for simultaneous equation solving ;)